OpenStudy (anonymous):

Solve the equation in the interval ( 0, 2pi ) tanx + secx= Sqrt 3

OpenStudy (danjs):

Recall: tan^2(x) + 1 = sec^2(x)

OpenStudy (michele_laino):

please note that your equation can be rewritten as below: \[\frac{ 1+\sin x }{ cos x }=\sqrt{3}\]

OpenStudy (anonymous):

I did get that far but I don't know what to do after that

OpenStudy (michele_laino):

so: \[1+\sin x=\sqrt{3}\cos x\] then: \[(1+\sin x)^{2}=3(\cos x)^{2}\] \[(1+\sin x)^{2}=3*(1-\sin x)(1+\sin x)\] finally: \[(1+\sin x)[[1+\sin x-3(1-\cos x)]=0\]

OpenStudy (anonymous):

oh ok

OpenStudy (michele_laino):

please semplify the term in square bracket, furthermore I used the fundamental identitu: cos^2+sin^2=1

OpenStudy (michele_laino):

oops: ...identity...

OpenStudy (michele_laino):

oops sorry: \[(1+\sin x)[1+\sin x-3(1-\sin x)]=0\]

OpenStudy (xapproachesinfinity):

hmm you don't need to square \(\sqrt{3}\cos x -\sin x =1 =====> \frac{\sqrt{3}}{2}\cos x-\frac{1}{2}\sin x=\frac{1}{2}\)

OpenStudy (anonymous):

how did you replace cosx^2 with (1-sinx)(1+sinx)

OpenStudy (xapproachesinfinity):

that's something we know already

OpenStudy (michele_laino):

yes, I did!

OpenStudy (anonymous):

xapproachesinfinity, i thought we already know that cos^2x= 1-sin^2x

OpenStudy (xapproachesinfinity):

well the way i did is much simpler an easier to get the answer + let me check the procedure @Michele_Laino did

OpenStudy (michele_laino):

more explanation: \[(1+\sin x)^{2}-3(1+\sin x)(1-\sin x)=0\] then I factor out (1+sin x): \[(1+\sin x)[(1+\sin x-3(1-\sin x)]=0\]

OpenStudy (xapproachesinfinity):

hmm seems fine with me you get 3 answers with that if my calculation are correct

OpenStudy (anonymous):

well now you two have figured it out and thanks for trying to help but i am still lost. I have a lot of other problems for my midterm study guide to do so thanks anyway

OpenStudy (michele_laino):

after a simple simplification, we have: \[4(1+\sin x)(\sin x-\frac{ 1 }{ 2 })=0\] so please you have to solve this subsequent equations: \[\sin x+1=0\] and: \[\sin x-\frac{ 1 }{ 2 }=0\]

OpenStudy (xapproachesinfinity):

in mine suggestion you have \(\cos(\pi/6)\cos x- \sin(\pi/6)\sin x=1/2==>\cos(\pi/6+x)=1/2 \)

OpenStudy (xapproachesinfinity):

not and! it is or :P

OpenStudy (xapproachesinfinity):

@0400782 did you understand what he did